a. List the sample space in the given experiment. How many outcomes are

possible?

b. Construct a table showing the number of defective computers in each

outcome and assign this number to this outcome. What is the value of the

random variable X?

c. Illustrate a probability distribution. What is the probability value P(X) to

each value of the random variable?

d. What is the sum of the probabilities of all values of the random

variable?

e. What do you notice about the probability of each value of the random

variable?


1
Expert's answer
2022-03-01T13:29:48-0500

a.There are "2^3=8" possible outcomes


"S=\{DDD, DDN, DND, NDD,"

"DNN,NDN, NND, NNN\}"


b. The possible values of the random variable "X" are "0, 1, 2, 3."

We will assume that the probability of getting heads and tails is the same:


"p = q =1\/2""\def\arraystretch{1.5}
 \begin{array}{c:c}
 Possible \ Outcomes & A \\ \hline
 DDD & 3 \\
 \hdashline
 DDN & 2 \\
 \hdashline
 DND & 2 \\
 \hdashline
 NDD & 2 \\
 \hdashline
 DNN& 1 \\
 \hdashline
 NDN & 1 \\
 \hdashline
 NND & 1 \\
 \hdashline
 NNN & 0 \\
 \hdashline
\end{array}"



c. Construct the probability distribution of the random variable


"\def\arraystretch{1.5}
 \begin{array}{c:c}
 x & 0 & 1 & 2 & 3 \\ \hline
 p(x) & 1/8 & 3/8 & 3/8 & 1/8 
\end{array}"

or


"\def\arraystretch{1.5}
 \begin{array}{c:c}
 x & 0 & 1 & 2 & 3 \\ \hline
 p(x) & 0.125 & 0.375 & 0.375 & 0.125 
\end{array}"



d.

"P(X=0)+P(X=1)+P(X=2)+P(X=3)"

"=1\/8+3\/8+3\/8+1\/8=1"

e.

The probability of each value of a discrete random variable is between 0 and 1, and the sum of all the probabilities is equal to 1.